# Exploring the AGN Fraction of a Sample of JWST’s Little Red Dots at : Overmassive Black Holes Are Strongly Favored

### v2

@@ 1 Introduction
JWST provides an unprecedented view of the early Universe. Unlike earlier telescopes that observed massive quasars at (Fan et al., 2001, 2003, 2006; Cool et al., 2006; Willott et al., 2009; Wang et al., 2007; Willott, 2011; Mortlock et al., 2011; Wu et al., 2015; Bañados et al., 2018; Fan et al., 2023), where the total spectral energy distribution (SED) is dominated by the emission of the central supermassive black hole (SMBH, with mass ), JWST can often capture the light from the host’s stellar component (see, e.g., Ding et al. 2023), especially in the case of lower-mass (i.e., ), lower-luminosity black holes (i.e., ). This extraordinary sensitivity has led to the detection of new populations of SMBHs in the high- Universe, farther and with lighter and more compact hosts (e.g., GN-z11 Maiolino et al. 2024a; Baggen et al. 2023; Greene et al. 2024).
@@
Efforts to quantify these scaling relations using JWST observations at have revealed a population of central SMBHs that are overmassive compared to the stellar content of their hosts (Maiolino et al., 2024a, 2023; Harikane et al., 2023; Larson et al., 2023; Übler et al., 2023; Stone et al., 2023; Furtak et al., 2024; Kocevski et al., 2023; Kokorev et al., 2023; Yue et al., 2024b; Juodžbalis et al., 2024; Kocevski et al., 2024), but seem to agree with the stellar velocity dispersion and the dynamical mass (Maiolino et al., 2023). More specifically, the relation inferred from JWST detections deviates by from the local relationship: black holes are overmassive by factors of to with respect to local counterparts hosted by a galaxy of the same size (Pacucci et al., 2023).
@@
However, determining the location of these high- systems in the plane crucially depends on how the contributions from the stellar and the active galactic nucleus (AGN) components of the SEDs are evaluated, as well as the individual uncertainties in the determination of the stellar mass and the black hole mass. This study investigates how different assumptions on the AGN/star ratio (i.e., how much these two components generate the observed flux) affect the goodness-of-fit with the observed photometry and the system’s location on the plane.
In particular, we analyze eight extremely red objects (EROs) first reported in Barro et al. 2024 and observed with JWST. As detailed in Sec. 2, these EROs are a subset of the more general population of “Little Red Dots”, or LRDs, a name that was first introduced by Matthee et al. (2023) and that is now of widespread use. Since the dataset from Barro et al. (2024) was published, several other studies reported the detection of similarly red and compact objects, which are ubiquitous in many JWST fields (see, e.g., the extensive catalog provided by Kocevski et al. 2024). In particular, Greene et al. (2024) argued that most (i.e., ) of these sources host a SMBH at their center, based on the detection of broad line regions (BLRs).
@@ 2 Data
Pérez-González et al. (2024) reports the discovery of 31 LRDs, whose selection is based on NIRcam colors F277W-F444W 1 mag, F150W-F200W 0.5 mag, and F444W 28 mag. These objects are selected from the JWST Advanced Deep Extragalactic Survey (JADES, Eisenstein et al. 2023), with matched Mid-IR observation from the SMILES survey. We select a sub-sample of 19 sources, which are characterized by MIRI detection, within and . We only adopt fluxes in the NIRCam filters (F090W, F115W, F150W, F182M, F200W, F210M, F277W, F335M, F356W, F410M, F430M, F444W, F460M, F480M), and MIRI filters (F560W, F770W, F1000W, F1280W, F1500W, F1800W, F2100W, F2550W). Similar to Pérez-González et al. 2024, we impose a 5 upper limit for filters where there are non-detections. We do not use any HST fluxes or upper limits in our analysis.
@@ 3 Methods: SED modeling
We fit the available photometry of the eight LRDs using SEDs built on a combination of two emission templates: stellar and AGN. We selected templates from the ones provided in CIGALE to vary only a small subset of parameters to focus our analysis on how the AGN fraction affects the SED fitting results. Table 1 lists the selected CIGALE templates and the parameters that were kept free, alongside the parameters where we adopted a different value from the default value. We note that we adopted the fiducial values of redshift for each source, as provided by Pérez-González et al. (2024); hence, CIGALE does not fit redshifts.
@@
For dust attenuation, we employ the dustatt_modified_starburst module based on the Calzetti et al. 2000 attenuation law, extended to lower wavelengths with Leitherer et al. 2002. The Ly result from Lin et al. 2024 employed a Small Magellanic Cloud dust law, yielding an E(B-V) color excess of 0.1, which we also adopt for our analysis. We employ the dl2014 module Draine et al. (2014) for galactic dust emission. For spatially unresolved sources at high redshift, there is little information regarding the relative contributions of the torus and the interstellar medium to the total dust attenuation (Hickox & Alexander, 2018). For our analysis, we do not vary dust attenuation on a galactic scale; instead, we choose to focus on the torus for simplicity.
@@ 4.1 AGN nature of LRD
This analysis aims to constrain the nature of the LRDs while accounting for our lack of information on precisely how much the AGN contributes to the observed SED. To this end, we model our photometry using both Type I unobscured AGN and Type II obscured AGN templates. For all 19 galaxies in our analysis, there is a clear preference for the Type I AGN template, as it consistently returned better reduced s than the obscured Type II template. The Type II templates typically yielded reduced values an order of magnitude larger than the Type I fits, often with such poor statistics that CIGALE could not return a best-fit model. Therefore, the remainder of this paper will focus on the results of the SED fitting using the Type I unobscured AGN template.
In addition to this, the reddening levels that are recovered from the Type I templates are more in agreement with the current landscape of LRD, where a significant amount is observed to have broad line emission (Greene et al., 2024).
@@ 4.2 Stellar-Dominated or AGN-Dominated?
The direct outputs from the SED fitting process outlined in Section 3 are the stellar mass of the host and the AGN bolometric luminosity. This latter parameter is then used to derive the mass of the central black hole by assuming that accretion occurs at the Eddington luminosity. Note that most of the lower-luminosity broad line AGN found at by JWST thus far have a high mean Eddington ratio of close to unity (see the related discussion in, e.g., Harikane et al. 2023; Maiolino et al. 2023; Pacucci & Loeb 2024). It is important to remark that lower values of the Eddington ratios would require a larger black hole mass, which would render our black holes even more overmassive. In principle, super-Eddington accretion rates may be possible in these high- environments given the wide availability of cold gas (Power et al., 2010; Pacucci & Loeb, 2020). However, for the purposes of this analysis, we assume a maximum luminosity of approximately , as has been observed in other AGN populations across a range of redshifts (e.g., Aird et al. 2012; Fan et al. 2023; Li et al. 2024).
@@ 4.3 Location in the Plane
Furthermore, we highlight that the presence of a black hole in these galaxies is supported by considerations on the limited amount of stellar mass that these hosts can contain, as they are very compact in size (see, e.g., Pacucci & Loeb 2024). In fact, typical effective radii are in the order of pc (Baggen et al., 2023). SED fitting executed with only stellar components (i.e., an AGN fraction of zero) would lead, in some cases, to average stellar densities that are larger, by factors , than core densities in globular clusters (Pacucci et al., in prep.).
@@
Figure 3a shows two galaxies for which the best fit AGN fraction is negligible (fracAGN = 2%), those two galaxies are Jades-187025 and Jades-204851. Jades-204851 as mentioned in 2, is the only broad line detected galaxy in our sample, and so it is concerning that it is only best fit by an AGN contribution of 2%. The surprisingly low best fit AGN contribution was suggestive of the two galaxies being star forming galaxies, instead of high redshift AGN hosts. For JADES-187025, this was indeed the case, as the SED fitting results for the stellar model returns a better reduced compared to the composite (AGN + stellar) model, and so we conclude that it is a dusty star forming galaxy. However, when we repeat the same analysis for JADES-204851, we find that the reduced for the stellar model is worse compared to the composite model.
Interestingly, Pérez-González et al. 2024 also recovered a low AGN contribution of approximately 10% when using a composite template of an unobscured QSO and a stellar template to fit the photometry for JADES-208451. Closer inspection of the CIGALE stellar model spectrum for JADES-204851 suggests that CIGALE could be underestimating the dust emission, which if correctly modeled could potentially lead to a much better SED fit for a stellar model. CIGALE modeling, much like the majority of the publicly available SED fitting codes, is based on an energy balance principle, where the absorbed light in the UV is remitted by the dust in the IR, this approach completely omits potential contributions from a hot dust component and has been found to be less reliable when fitting star forming galaxies at high redshifts (z 4, Haskell et al. 2023). And so it is plausible that JADES-204851 could be a high redshift star forming galaxy, whose broad line emission is due to kinematics and not AGN activity (Baggen et al., 2024).
@@ 4.3.1 AGN contribution at the Balmer limit, H∞
The spectral slope break at rest H∞ discussed in Setton et al. 2024 implies that an AGN model is inconsistent with explaining the spectroscopy of LRD. Here we asses the strength of the AGN contribution at the Balmer limit to determine the consistency of the composite model presented here under consideration of the spectral slope break. To constrain the AGN contribution at the Balmer limit, we define a new AGN fraction at rest H∞ (3645 Å). We then define a threshold of 50% to represent where the AGN contribution would impact the spectral slope at rest H∞. Figure 3b shows the location along the where the AGN contribution at the Balmer limit exceed 50 percent. We find that for most of the galaxies analyzed here (16/19) an AGN contribution that exceed 50% at the Balmer limit is consistent with a ( 90%) mid-IR AGN fraction. Such a high mid-IR AGN fraction is well beyond the typical ranges we recover as best fit AGN fractions, even in cases where the black hole masses are overmassive by up to a factor of 4, see Figure 3a. Therefore, we conclude that for the population of LRD with spectral slope break at rest H∞, a significant mid-IR AGN ( 70%) component that can produce an overmassive black hole result would preserve the spectral feature at rest H∞.
Figure 4 shows the continuum from the CIGALE model for the seven galaxies with F277W - F444W 1.7. For three out of the seven galaxies, there is a prominent spectral break at the Balmer limit, similar to what was reported by Setton et al. 2024. All three galaxies feature a best fitting AGN fraction in the rest-frame mid-IR between 40 and 70 percent. This shows that a reddened AGN component can account for the extra red emission in the optical for LRD while maintaining the H∞ spectral break. We must emphasize however, that this result does not rule out other possible sources for the optical red emission that is observed in LRD.
@@ 4.3.2 Photometric selection of sources
The LRD selection criteria adopted by Pérez-González et al. 2024 is largely consistent with other works (see Greene et al. 2024; Kocevski et al. 2023; Barro et al. 2024 ) in terms of color and compactness. The only discernible difference is the F277W - F444W magnitude criteria, where Pérez-González et al. 2024 uses a threshold of 1, other works used a threshold of 1.7. To test how this criteria affects our analysis, we restrict our sample to only include galaxies where the F277W - F444W is greater than 1.7. Adopting this color criteria reduced our galaxy catalog from nineteen to seven. Figure 3c shows the parameter space for the seven reddest galaxy in our catalog. Despite the smaller sample size, we find that our results still holds – the mid-IR AGN contribution needed to produce overmassive black holes is still significant ( 40%), while the threshold for the AGN emission to affect the spectral slope at rest H∞ is typically much higher than the best fit mid-IR AGN fraction.
@@ 4.4.1 MIRI photometry
We also found that the depth of the included MIRI photometry affects the precision of CIGALE estimates of stellar masses. Specifically, we find that as the AGN contribution becomes maximal, the precision of the CIGALE stellar mass estimate worsens. However, we find that the precision becomes artificially enhanced as the included MIRI wavelengths becomes shorter. This result suggests that SED fitting analysis of LRD where the only considered AGN fraction is maximal could have some under-estimated errors on their stellar mass estimates, especially in cases where the MIRI photometry is not included.
@@ 4.4.2 Polar reddening of AGN torus
In addition to mid-IR photometry, we also found that the recovered stellar mass and AGN luminosity were highly sensitive to how we modeled the polar reddening, . Here we discuss how the modeling choices for the polar reddening, such as the attenuation law, and the flexibility of the reddening parameter,E(B-V), affected our analysis results.
Recent spectroscopic analysis of LRD (Greene et al., 2024; Maiolino et al., 2023) have commonly modeled the AGN polar reddening using a Small Magellanic Cloud (SMC) dust law from Pei 1993. The attenuation law used in our analysis follows fro the Balmer decrement analysis of LRDs from Brooks et al. 2024, as mentioned in section 3. To compare the effects of attenuation law on the SED fitting results, we re-ran our analysis using a SMC dust law. Using the SMC dust law resulted in a ( 10%) higher AGN luminosity compared to using the Calzetti et al. 2000 dust law.
In addition to the choice of attenuation law, we also tested how the model choice for the reddening parameter would affect our analysis. We initially allowed CIGALE to sample values between 1 and 4, corresponding to an range of 3.63 to 12, assuming a (Calzetti et al., 2000) extinction law. This sampling range is consistent with recent Balmer decrement analyses, which indicated that the ratio in LRDs suggests AGN reddening with (Brooks et al., 2024).
However, we found that the recovered parameter space performed poorly, as shown in Figure 6, especially at lower AGN fractions, where the reddening was typically higher, ranging between 1.4 and 3. Despite this, for all 19 galaxies in our sample, the best-fit AGN fraction occurred consistently at an value of 1. Motivated by this behavior, we re-ran our analysis with a fixed value of 1. As shown in Figure 5, this fixed polar AGN reddening behaves more smoothly across the parameter space compared to the results in Figure 6.

### v3

@@ 1 Introduction
JWST provides an unprecedented view of the early Universe. Unlike earlier telescopes that observed massive quasars at (Fan et al., 2001, 2003, 2006; Cool et al., 2006; Willott et al., 2009; Wang et al., 2007; Willott, 2011; Mortlock et al., 2011; Wu et al., 2015; Bañados et al., 2018; Pacucci & Loeb, 2019; Fan et al., 2023), where the total spectral energy distribution (SED) is dominated by the emission of the central supermassive black hole (SMBH, with mass ), JWST can often capture the light from the host’s stellar component (see, e.g., Ding et al. 2023), especially in the case of lower-mass (i.e., ), lower-luminosity black holes (i.e., ). This extraordinary sensitivity has led to the detection of new populations of SMBHs in the high- Universe, farther and with lighter and more compact hosts (e.g., GN-z11 Maiolino et al. 2024a; Baggen et al. 2023; Greene et al. 2024).
@@
Efforts to quantify these scaling relations using JWST observations at have revealed a population of central SMBHs that are overmassive compared to the stellar content of their hosts (Maiolino et al., 2024a, b; Harikane et al., 2023; Larson et al., 2023; Übler et al., 2023; Stone et al., 2023; Furtak et al., 2024; Kocevski et al., 2023; Kokorev et al., 2023; Yue et al., 2024b; Juodžbalis et al., 2024; Kocevski et al., 2024; Taylor et al., 2024), but seem to agree with the stellar velocity dispersion and the dynamical mass (Maiolino et al., 2024b). More specifically, the relation inferred from JWST detections deviates by from the local relationship: black holes are overmassive by factors of to with respect to local counterparts hosted by a galaxy of the same size (Pacucci et al., 2023).
@@
However, determining the location of these high- systems in the plane crucially depends on how the contributions from the stellar and the active galactic nucleus (AGN) components of the SEDs are evaluated, as well as the individual uncertainties in the determination of the stellar mass and the black hole mass (in this regard, see also, e.g., Pacucci et al. 2022). This study investigates how different assumptions on the AGN/star ratio (i.e., how much these two components generate the observed flux) affect the goodness-of-fit with the observed photometry and the system’s location on the plane.
In particular, we analyze 19 little red dots (LRDs) first reported in Barro et al. 2024 and observed with JWST. As detailed in Sec. 2, these EROs are a subset of the more general population of “Little Red Dots”, or LRDs, a name that was first introduced by Matthee et al. (2024) and that is now of widespread use. Since the dataset from Barro et al. (2024) was published, several other studies reported the detection of similarly red and compact objects, which are ubiquitous in many JWST fields (see, e.g., the extensive catalog provided by Kocevski et al. 2024). In particular, Greene et al. (2024) argued that most (i.e., ) of these sources host a SMBH at their center, based on the detection of broad line regions (BLRs).
@@ 2 Data
Pérez-González et al. (2024) reports the discovery of 31 LRDs, whose selection is based on NIRcam colors F277W-F444W 1 mag, F150W-F200W 0.5 mag, and F444W 28 mag. These objects are selected from the JWST Advanced Deep Extragalactic Survey (JADES, Eisenstein et al. 2023), with matched Mid-IR observation from the SMILES survey. We select a sub-sample of 19 sources, which are characterized by MIRI detection, within and . We only adopt fluxes in the NIRCam filters (F090W, F115W, F150W, F182M, F200W, F210M, F277W, F335M, F356W, F410M, F430M, F444W, F460M, F480M), and MIRI filters (F560W, F770W, F1000W, F1280W, F1500W, F1800W, F2100W, F2550W). Similar to Pérez-González et al. 2024, we impose a 5 upper limit for filters where there are non-detections. For the galaxies with non-detections in the short NIRCam wavelengths, we supplement the photometry using 0.25” aperture HST fluxes and 5 upper limits in both F606W and F814W, as provided in the JADES catalog.
@@ 3 Methods: SED modeling
We fit the available photometry of the eight LRDs using SEDs built on a combination of two emission templates: stellar and AGN. We selected templates from the ones provided in CIGALE to vary only a small subset of parameters to focus our analysis on how the AGN fraction affects the SED fitting results. Table 1 lists the selected CIGALE templates and the parameters that were kept free, alongside the parameters where we adopted a different value from the default value. We note that we adopted the fiducial redshift values for each source, as provided by Pérez-González et al. (2024); hence, CIGALE does not fit redshifts.
@@
For dust attenuation, we employ the dustatt_modified_starburst module based on the Calzetti et al. 2000 attenuation law, extended to lower wavelengths with Leitherer et al. 2002. The Ly result from Lin et al. 2024 employed a Small Magellanic Cloud dust law, yielding an E(B-V) color excess of 0.1, which we also adopt for our analysis. We employ the dl2014 module Draine et al. (2014) for galactic dust emission. For spatially unresolved sources at high redshift, there is little information regarding the relative contributions of the torus and the interstellar medium to the total dust attenuation (Hickox & Alexander, 2018). For our analysis, we do not vary dust attenuation on a galactic scale; instead, we focus on the torus for simplicity.
@@ 4.1 AGN nature of LRD
This analysis aims to constrain the nature of the LRDs while accounting for our lack of information on precisely how much the AGN contributes to the observed SED. To this end, we model our photometry using both Type I unobscured AGN and Type II obscured AGN templates. For all 19 galaxies in our analysis, there is a clear preference for the Type I AGN template, as it consistently returned better reduced s than the obscured Type II template. The Type II templates typically yielded reduced values an order of magnitude larger than the Type I fits, often with such poor statistics that CIGALE could not return a best-fit model. Additionally, for galaxies where the AGN component is consistent with a type II template, AGN contribution to the rest optical wavelength emission is negligible, and so we do not explore these any further. Therefore, the remainder of this paper will focus on the results of the SED fitting using the Type I unobscured AGN template.
In addition to this, we found the recovered parameter space showed behavior that appears not to be physically motivated, such as little to no evolution in the estimated black hole mass as the AGN fraction increases, smaller AGN fraction corresponding to a larger black hole mass compared to higher AGN fractions in the same galaxy. We found the CIGALE code sensitive to the torus reddening parameter. Thus, our choice to allow the parameter to vary from 1 to 4 resulted in some extreme modeling configurations. However, we find that for all 19 galaxies, the best-fit AGN fraction occurred consistently at an value of 1. Therefore, we fix the torus reddening to a value of 1 and re-ran our analysis, the results of which will be discussed further below.
@@ 4.2 Stellar-Dominated or AGN-Dominated?
The direct outputs from the SED fitting process outlined in Section 3 are the stellar mass of the host and the AGN bolometric luminosity. This latter parameter is then used to derive the mass of the central black hole by assuming that accretion occurs at the Eddington luminosity. Note that most of the lower-luminosity broad line AGN found at by JWST thus far have a high mean Eddington ratio of close to unity (see the related discussion in, e.g., Harikane et al. 2023; Maiolino et al. 2024b; Pacucci & Loeb 2024). It is important to remark that lower values of the Eddington ratios would require a larger black hole mass, which would render our black holes even more overmassive. In principle, super-Eddington accretion rates may be possible in these high- environments given the wide availability of cold gas (Power et al., 2010; Pacucci & Loeb, 2020). However, for this analysis, we assume a maximum luminosity of approximately , as has been observed in other AGN populations across a range of redshifts (e.g., Aird et al. 2012; Fan et al. 2023; Li et al. 2024).
@@ 4.3 Location in the Plane
Furthermore, we highlight that the presence of a black hole in these galaxies is supported by considerations of the limited amount of stellar mass that these hosts can contain, as they are very compact in size (see, e.g., Pacucci & Loeb 2024). In fact, typical effective radii are in the order of pc (Baggen et al., 2023). SED fitting executed with only stellar components (i.e., an AGN fraction of zero) would lead, in some cases, to average stellar densities that are larger, by factors , than core densities in globular clusters (Pacucci et al., in prep.).
@@
Figure 3a shows two galaxies for which the best fit AGN fraction is negligible (fracAGN = 2%), those two galaxies are Jades-187025 and Jades-204851. Jades-204851 as mentioned in 2, is the only broad line detected galaxy in our sample, and so it is concerning that it is only best fit by an AGN contribution of 2%. The surprisingly low best-fit AGN contribution was suggestive of the two galaxies being star-forming galaxies, instead of high redshift AGN hosts. For JADES-187025, this was indeed the case, as the SED fitting results for the stellar model return a better reduced compared to the composite (AGN + stellar) model, and so we conclude that it is a dusty star-forming galaxy. However, when we repeat the same analysis for JADES-204851, we find that the reduced for the stellar model is worse compared to the composite model.
Interestingly, Pérez-González et al. 2024 also recovered a low AGN contribution of approximately 10% when using a composite template of an unobscured QSO and a stellar template to fit the photometry for JADES-208451. Closer inspection of the CIGALE stellar model spectrum for JADES-204851 suggests that CIGALE could be underestimating the dust emission, which, if correctly modeled, could potentially lead to a much better SED fit for a stellar model. CIGALE modeling, much like the majority of the publicly available SED fitting codes, is based on an energy balance principle, where the absorbed light in the UV is remitted by the dust in the IR; this approach completely omits potential contributions from a hot dust component and is less reliable when fitting star-forming galaxies at high redshifts (z 4, Haskell et al. 2023). And so it is plausible that JADES-204851 could be a high redshift star-forming galaxy, whose broad line emission is due to kinematics and not AGN activity (Baggen et al., 2024).
@@ 4.3.1 AGN contribution at the Balmer limit, H∞
The spectral slope break at rest H∞ discussed in Setton et al. 2024 implies that an AGN model is inconsistent with explaining the spectroscopy of LRD. Here we assess the strength of the AGN contribution at the Balmer limit to determine the consistency of the composite model presented here under consideration of the spectral slope break. To constrain the AGN contribution at the Balmer limit, we define a new AGN fraction at rest H∞ (3645 Å). We then define a threshold of 50% to represent where the AGN contribution would impact the spectral slope at rest H∞. Figure 3b shows the location along the where the AGN contribution at the Balmer limit exceeds 50 percent. We find that for most of the galaxies analyzed here (16/19) an AGN contribution that exceed 50% at the Balmer limit is consistent with a ( 90%) mid-IR AGN fraction. Such a high mid-IR AGN fraction is well beyond the typical ranges we recover as best-fit AGN fractions, even in cases where the black hole masses are overmassive by up to a factor of 4, see Figure 3a. Therefore, we conclude that for the population of LRD with spectral slope break at rest H∞, a significant mid-IR AGN ( 70%) component that can produce an overmassive black hole result would preserve the spectral feature at rest H∞.
Figure 4 shows the continuum from the CIGALE model for the seven galaxies with F277W - F444W 1.7. For three out of the seven galaxies, there is a prominent spectral break at the Balmer limit, similar to what was reported by Setton et al. 2024. All three galaxies feature a best-fitting AGN fraction in the rest-frame mid-IR between 40 and 70 percent. This shows that a reddened AGN component can account for the extra red emission in the optical for LRD while maintaining the H∞ spectral break. We must emphasize, however, that this result does not rule out other possible sources for the optical red emission that is observed in LRD.
@@ 4.3.2 Photometric selection of sources
The LRD selection criteria adopted by Pérez-González et al. 2024 is largely consistent with other works (see Greene et al. 2024; Kocevski et al. 2023; Barro et al. 2024 ) in terms of color and compactness. The only discernible difference is the F277W - F444W magnitude criteria, where Pérez-González et al. 2024 uses a threshold of 1, other works used a threshold of 1.7. To test how this criterion affects our analysis, we restrict our sample only to include galaxies where the F277W - F444W is greater than 1.7. Adopting this color criterion reduced our galaxy catalog from nineteen to seven. Figure 3c shows the parameter space for the seven reddest galaxies in our catalog, while Table 2 reports their redshifts.
Source ID | Redshift |
81400 | 5.9 |
121710 | 7.1 |
132229 | 7.247 |
154428 | 6.251 |
203749 | 4.1153 |
204851 | 5.48 |
219000 | 6.8119 |
Despite the smaller sample size, we find that our results still hold – the mid-IR AGN contribution needed to produce overmassive black holes is still significant ( 40%), while the threshold for the AGN emission to affect the spectral slope at rest H∞ is typically much higher than the best-fit mid-IR AGN fraction.
@@ 4.4.1 MIRI photometry
We also found that the depth of the included MIRI photometry affects the precision of CIGALE estimates of stellar masses. Specifically, we find that as the AGN contribution becomes maximal, the precision of the CIGALE stellar mass estimate worsens. However, we find that the precision becomes artificially enhanced as the included MIRI wavelengths become shorter. This result suggests that SED fitting analysis of LRD where the only considered AGN fraction is maximal could have some under-estimated errors on their stellar mass estimates, especially in cases where the MIRI photometry is not included.
The relevance of MIRI detections in our sample of LRDs warrants further discussion. Previous studies of LRDs (see, e.g., Williams et al. 2024; Akins et al. 2024) remarkably found very faint MIRI fluxes in the F770W and F1800W filters. Overall, this finding suggests a lack of hot dust emission. However, Li et al. (2025) demonstrated that this faint emission in MIRI bands can be easily explained by an extended density distribution in the dust torus model. Other studies (see, e.g., Wang et al. 2024) highlighted the presence of Balmer breaks as indicator of the prominence of stellar activity (however, see also Inayoshi & Maiolino 2025), and they also suggested that deeper and redder data are required to investigate the nature of the LRDs. The fact that this study focuses on LRDs with strong MIRI detections inherently limits our analysis to systems characterized by bright dust emission in the infrared.
@@ 4.4.2 Polar reddening of AGN torus
In addition to mid-IR photometry, we also found that the recovered stellar mass and AGN luminosity were highly sensitive to how we modeled the polar reddening, . Here, we discuss how the modeling choices for the polar reddening, such as the attenuation law and the flexibility of the reddening parameter, E(B-V), affected our analysis results.
Recent spectroscopic analysis of LRDs (Greene et al., 2024; Maiolino et al., 2024b) have commonly modeled the AGN polar reddening using a Small Magellanic Cloud (SMC) dust law from Pei 1993. The attenuation law used in our analysis follows from the Balmer decrement analysis of LRDs from Brooks et al. 2024, as mentioned in section 3. To compare the effects of attenuation law on the SED fitting results, we re-ran our analysis using a SMC dust law. Using the SMC dust law resulted in a ( 10%) lower AGN luminosity compared to using the Calzetti et al. 2000 dust law.
In addition to the choice of attenuation law, we also tested how the model choice for the reddening parameter would affect our analysis. Similar to our modeling with the Calzetti et al. 2000 dust law, we initially allowed CIGALE to sample values between 1 and 4. We find that even for the SMC dust law, the parameter space performed poorly, and only when the torus reddening is fixed to a value of 1 do we then get a monotonically behaved parameter space. This result suggests that the estimated stellar masses and black hole masses from photometric modeling of LRD are more sensitive to how the torus reddening is modeled and less the specific dust law used for modeling.
Figure 5 shows the space for all 19 galaxies modeled assuming a SMC dust law for the torus dust reddening, where the E(B-V)polar was sampled from 1 to 4. The behavior of the parameter space appears to be random, where, in some instances, the black hole mass at a lower AGN fraction is higher than at a higher AGN fraction. Additionally, at lower AGN fractions, we found the reddening to be systematically higher, ranging between 1.4 and 3, which would be inconsistent with any Hα emission being detected in these galaxies. Fixing the polar reddening value to 1 results in the parameter space shown in Figure 6, where the relationship between the AGN fraction and the black hole mass is as expected.